qboy:

A chocolate bar measures 50 mm wide, 90 mm long, and 3 and 1 over 2 mm high. Will’s Wrapping Company is making a wrapper to cover the chocolate bar. Use the correct net and determine how much paper will be needed to make the wrapper. The net of a square prism with a height of 50 millimeters, a length of 90 millimeters, and a width of 3 and 1 over 2 millimeters.

generation2006:

A=2(xy+xz+yz, where x,y and z are the three dimensions of the prism A=2(50*90+50*3.5+90*3.5) A=2(4500+175+315) A=2(4990) which means A is 2 times 4990 and at the end of the result just add mm^2

generation2006:

which would look like ____mm^2 (the blank like this _____represent whatever you get as you result)

generation2006:

you result of 2 times 4900